OpenStudy (anonymous):

The cost of three notebooks and four pencils is $8.50. The cost of five notebooks and eight pencils is $14.5". Determine the cost of one notebook and the cost of one pencil. (Needs an algebraic solution)

OpenStudy (anonymous):

$14.50*

OpenStudy (anonymous):

lets call x notebooks and y pencils

OpenStudy (anonymous):

\[3N +4P = 8.50\]\[5N + 8P = 14.50\]Multiply the top equation by -2. \[-6N - 8P = -17\] use elimination to solve. -6N-8P = -17 5N +8P = 14.50 Result: -N = -2.50 N(otebook) = 2.50 Plug in N to either equation. 3(2.50) +4p = 8.50 4P + 7.5 = 8.50 4P = 1 P = .25

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

the first equation will then look like 3x + 4y =$8.50